# EnsemW2S: Can an Ensemble of LLMs be Leveraged to Obtain a Stronger LLM?

### v1

@@ Abstract
How can we harness the collective capabilities of multiple Large Language Models (LLMs) to create an even more powerful model? This question forms the foundation of our research, where we propose an innovative approach to weak-to-strong (w2s) generalization—a critical problem in AI alignment. Our work introduces an easy-to-hard (e2h) framework for studying the feasibility of w2s generalization, where weak models trained on simpler tasks collaboratively supervise stronger models on more complex tasks. This setup mirrors real-world challenges, where direct human supervision is limited. To achieve this, we develop a novel AdaBoost-inspired ensemble method, demonstrating that an ensemble of weak supervisors can enhance the performance of stronger LLMs across classification and generative tasks on difficult QA datasets. In several cases, our ensemble approach matches the performance of models trained on ground-truth data, establishing a new benchmark for w2s generalization. We observe an improvement of up to 14% over existing baselines and average improvements of 5% and 4% for binary classification and generative tasks, respectively. This research points to a promising direction for enhancing AI through collective supervision, especially in scenarios where labeled data is sparse or insufficient.
@@ 1 Introduction
As AI models, particularly Large Language Models (LLMs), continue to surpass human performance in various domains, a pressing challenge arises: how do we effectively supervise models that exceed our capabilities? This problem, known as super-alignment, is exacerbated by the scarcity of high-quality labeled data, which limits direct human oversight. The key question driving our work is whether weak models, trained on simpler tasks, can be leveraged to instruct and improve stronger models in complex settings—a problem known as weak-to-strong (w2s) generalization.
The concept of w2s generalization was introduced by Burns et al. (2023), where weak models are used to align stronger models in the absence of sufficient ground-truth supervision. However, while this work laid the groundwork, it left several critical challenges unresolved. (C1) Single Weak Supervisor Limitation. Prior studies (Burns et al., 2023; Ji et al., 2024; Charikar et al., 2024; Lang et al., 2024) tend to rely on a single weak supervisor, limiting the diversity and robustness of the supervision. A single model’s perspective often falls short when attempting to instruct stronger models in more complex tasks, highlighting the need for a more diversified supervisory approach. (C2) Lack of Focus on Weak Model Enhancement. Another limitation is that previous research (Burns et al., 2023; Ji et al., 2024; Charikar et al., 2024; Lang et al., 2024) has focused predominantly on improving knowledge transfer from weak to strong models without addressing how to enhance the weak models themselves. This oversight leaves weak models under-optimized, thereby restricting their utility in complex problem settings. (C3) Overlooking Task Complexity. Furthermore, while task complexity plays a crucial role in determining how well weak models can supervise stronger ones, most prior work (Sun et al., 2024) has not adequately addressed this issue. For instance, Burns et al. (2023) briefly explored the impact of task complexity using chess data, but a more structured and systematic approach is needed to differentiate between easy and hard tasks and study their effects on supervision.
To address these challenges, we propose a novel ensemble-based method designed to improve w2s generalization. Central to our approach is an easy-to-hard (e2h) framework, which extends w2s generalization by focusing on the progression from simpler tasks (easy) to more complex tasks (hard). This mirrors practical scenarios, where human oversight is more feasible for simpler tasks, and weak models must step in to guide stronger models in tackling harder tasks. In this setting, weak models trained on easy data supervise stronger models working on more difficult problems, creating a more pragmatic approach to w2s generalization.
To further enhance the capabilities of weak models, we develop a novel AdaBoost-inspired ensemble method for generation tasks, in addition to classification tasks. By combining the supervision of multiple weak models, we create a more robust and effective supervisory system for stronger LLMs. This ensemble approach overcomes the limitations of single-supervisor systems and introduces a mechanism to refine the weak models themselves, ensuring they can provide meaningful guidance even in complex tasks. Our experiments demonstrate that this ensemble method not only improves the weak models’ generalization capabilities but also enables stronger models to achieve performance on par with oracle models trained on high-quality data.
@@
(1) We introduce an ensemble method inspired by AdaBoost, combining weak LLMs to provide stronger supervision for training stronger models. Our approach is validated through experiments on binary classification tasks, where we observe improvements of up to 14% over baselines and an average improvement of 7% across all model pairs, showcasing the feasibility of w2s generalization.
(2) We extend this framework to supervised fine-tuning tasks for autoregressive LLMs, where our novel algorithm combines weak LLMs via a voting mechanism that adjusts token probabilities. In several cases, we observe our strong model trained using weak labels to outperform the strong model trained on ground truth, thus enabling effective supervision, even on complex tasks.
(3) We propose a practical easy-to-hard (e2h) framework for w2s generalization, where models trained on easy data provide supervision for harder tasks. This setup emphasizes the importance of task complexity and demonstrates significant improvements when weak models guide strong LLMs. For our EnsemW2S-AdaBoost method, along with observing w2s-trained student models outperforming the strong student oracle in several e2h generalization scenarios, we also observe accuracy improvements of up to 10% over baselines and an average improvement of 3.34% and 4.4% for Quartz and ARC data respectively.
2 Weak-to-Strong Generalization via Easy-to-Hard Framework
Figure 1: This figure illustrates the complete pipeline of our EnsemW2S method for easy-to-hard generalization using w2s generalization. In a realistic scenario, weak experts are adept at answering easy questions but must supervise strong models to tackle hard problems. In the leftmost portion, we show that we train weak models on easy data, strong models on hard data, and transfer models on pseudo labels generated by the weak model on hard data. Ultimately, we aim to increase the Performance Gap Recovered (PGR). On the right, we depict how our EnsemW2S-AdaBoost algorithm chooses the correct answer at the token level. At the bottom, we provide an example of easy and hard data for the Quartz dataset for e2h generalization, highlighting the importance of distinguishing between easy and hard data for realistic w2s generation.
The Overall Idea. We investigate the easy-to-hard framework as a more pragmatic setting to study the (im)possibility of w2s generalization. In this framework, weak models train on simpler tasks and subsequently instruct strong models to tackle more complex challenges, closely mirroring real-world conditions with limited human oversight. Figure 1 explains our idea and pipeline for easy-to-hard generalization using w2s generalization. (Figure 7 in the Appendix provides the detailed algorithmic and data flow). In a realistic scenario, weak experts are proficient in answering easy questions but must supervise strong models to tackle hard problems. We train weak models on easy data and strong models on hard data. A transfer model is trained using pseudo labels generated by the weak model on the hard data. Ultimately, we aim to improve the Performance Gap Recovered (PGR).
2.1 The Easy-to-Hard Framework
Weak Model $h_{\theta}$ as the Teacher. A state-of-the-art LLM $h_{\theta}$ is trained on a set of ‘easy data’ that we currently have access to labels, i.e., $(\bm{x}^{e},\bm{y}^{e})$. For example, this could be Go games, math problems, or common sense reasoning questions that we have solutions for. This ‘weak teacher’ is trained on the labeled easy data $(\bm{x}^{e},\bm{y}^{e})$. Although we refer to this model as a “weak teacher”, it is only relatively weak compared to the strong model we aim to obtain. Moreover, the “easy data” is only relatively easy compared to the hard data for which we currently lack solutions. Thus, the easy data may not be simple but slightly easier than the hard data, which are currently unsolvable using existing models.
Strong Model $u_{\phi}$ as the Upper Bound. As an important part of our thought experiment, we establish an upper bound, which is not attainable in practice. Specifically, we assume access to the ground-truth labels of the hard data $(\bm{x}^{h},\bm{y}^{h})$, which is impractical but establishes an upper bound for this thought experiment. A model $u_{\phi}$, larger than the weak teacher $h_{\theta}$, is trained on the labeled hard data $(\bm{x}^{h},\bm{y}^{h})$. The reason why $u_{\phi}$ is larger than $h_{\theta}$ is that we believe a model strong enough to solve hard questions that no existing models can solve will require high capacity.
Weak-to-Strong Model $f_{\phi}$ Obtained in Practice. To test the weak-to-strong generalization, we will train a weak-to-strong transfer model $f_{\phi}$ that has the same capacity as the strong model, i.e., the same model size as $u_{\phi}$, but is not trained under the unrealistic assumption of oracle access to hard labels. Rather, it is trained using weak teacher’s feedback. Specifically, we consider using the pseudo-labeled $(\bm{x}^{h},h_{\theta}(\bm{x}^{h})$ as training data for training the weak-to-strong transfer model $f_{\phi}$.
2.2 Easy and Hard Data
Dataset and Setup. We use the SciQ dataset (Welbl et al., 2017) for the binary classification task. It is a multiple-choice science question-answer dataset and is also used as one of the NLP classification datasets by Burns et al. (2023). We convert it into binary labels following (Burns et al., 2023). For the supervised fine-tuning (SFT) task on the Q/A dataset, we use ARC (Clark et al., 2018) and Quartz (Tafjord et al., 2019) datasets, which are also multiple-choice question-answer datasets, allowing us to generate multiple-choice pseudo labels. Ding et al. (2024) provide difficulty levels for some common mathematics and programming problems, chess puzzles, and reasoning question datasets, which can be further utilized to expand this work.
Easy $(\bm{x}^{e},\bm{y}^{e})$ and Hard $(\bm{x}^{h},\bm{y}^{h})$ Data Split. To generate difficulty ratings for our datasets, we employ the $n$-fold cross-validation method. We train the model on the $(n-1)$ out of $n$ splits of the data and test on the remaining split. We repeat the process $n$ times with different splits for testing each time and aggregate the errors. We use this error value for each sample as its difficulty rating. We split the low difficulty-rated data for weak model training and use the high difficulty-rated data to generate strong model training data and testing data randomly. We follow the same cross-validation method, with different training protocols, for generating difficulty for both binary classification and generation tasks. More details and our difficulty rating plots can be seen in Figures 8, 9, and 10 in the Appendix.
2.3 An Ensemble of Teachers
In a practical situation, we may face a dearth of strong supervisors but have an abundance of weak supervisors. Previous works (Burns et al., 2023; Ji et al., 2024) have used only one weak supervisor. Our work aims to combine the power of multiple weak supervisors to provide stronger supervision for better weak-to-strong (w2s) generalization. However, combining multiple weak supervisors to improve w2s generalization is challenging. In the following section, we detail how to combine a collection of weak teachers with diverse skill sets to obtain a competitive weak-to-strong model that is better than the weak model and ideally reaches or even surpasses the strong model, i.e., the upper bound of performance.
3 W2S Generalization via AdaBoost of Existing SoTA Diverse Teacher LLMs
In this section, we introduce our method to boost experts for two tasks: a binary classification task for an NLP dataset and a supervised fine-tuning task for multiple-choice Q/A datasets. A list of important notions is mentioned in Appendix D for reference.
3.1 AdaBoost of Weak LLM Teachers for Classification Tasks
This simple thought experiment tests weak-to-strong (w2s) generalization and is the first task tested by Burns et al. (2023). We utilize the vanilla AdaBoost algorithm (Algorithm 2 detailed in the Appendix) to generate answers to a hard question $\bm{x}^{h}$ from each weak LLM teacher, i.e., generate $h_{\theta}^{t}(\bm{x}^{h})$ for $t\in\{1,\ldots,T\}$. A weighted “majority vote/aggregation” is implemented to generate a consensus as the answer $\mathbbm{1}(\sum_{t=1}^{T}\alpha_{t}h_{\theta}^{t}(\bm{x}^{h})>0)\in\{0,1\}$, also known as the pseudo-label, to the hard question $\bm{x}^{h}$. Here, the coefficients $\{\alpha_{t}\mid t\in\{1,\ldots,T\}\}$ are hyperparameters learned during the AdaBoost training.
AdaBoost utilizes the wisdom of the crowd to obtain a stronger learner. Inspired by its philosophy, we use an ensemble of weak LLM teachers as the “weak learners” to obtain a “stronger learner”, i.e., a strong model that improves binary classification tasks, thus achieving weak-to-strong generalization.
Results and Observations. As shown in the weak-model performance columns in Table 1 and 2 in the Appendix, the combined weak experts ($T=2,3,4,5$) demonstrate higher performance than a single weak expert (baseline).
Training Methodology. In our method, the weak experts are trained to minimize the error on the reweighted training examples, as detailed in Line 5 of Algorithm 2. The only requirement is that they perform better than random, thus satisfying the well-known weak learning condition. These weak experts represent a practical scenario where, although individually weak, they possess complementary knowledge. Thus, when combined, they have the potential to form a stronger expert.
3.2 Improving AdaBoost for Complex Generation Tasks
Challenges of Applying AdaBoost. The canonical AdaBoost algorithm assumes a sophisticated ensemble of feedback in the form of scores. However, LLMs are generative AI models known for their remarkable ability to generate coherent, free-form text. Applying the vanilla AdaBoost algorithm directly to generation tasks is challenging because (1) the output is not just a single class label but a sequence of text with no fixed length, and (2) different teachers may generate answers in various formats, making it non-trivial to combine their responses.
EnsemW2S-AdaBoost: Our modified AdaBoost Algorithm for Multiple-Choice Q/A Task. To address these challenges, we propose a modified multi-class AdaBoost algorithm where the number of classes corresponds to the vocabulary size. We treat each token as an independent sample, as shown in Algorithm 1, and apply multi-class AdaBoost (Hastie et al., 2009) with modifications, calling our algorithm EnsemW2S-AdaBoost.
@@ 3.2 Improving AdaBoost for Complex Generation Tasks
Token-Level Weighting. The first modification involves generating weights for each token within a sentence sample. We define the initial token-sample weights vector $D_{1}(i,j)\leftarrow\frac{1}{n}$ for all $i\in[m],j\in[k_{i}]$, where $n=\sum_{i=1}^{m}k_{i}$, $k_{i}$ is the number of tokens in the answer part of each sample $i$, $m$ is the total number of training data samples and $j$ is the $j^{\text{th}}$ token in a particular chosen $i^{\text{th}}$ sample. We update these weights, $D_{t}(i,j)$, for each iteration $t$ of EnsemW2S-AdaBoost.
@@
Token-Level Data Sampling. We sample $S^{\prime}=\{(\bm{x^{\prime}}^{e}_{i},\bm{y^{\prime}}^{e}_{i})\}_{i=1}^{m}$ from $S$ using token-sample weights $D_{t}(i,j)$. By sampling with respect to probability masses $D_{t}(i,j)$ with repetition, we obtain a set of $n=\sum_{i=1}^{m}k_{i}$ tokens to train on. However, treating these $n$ sampled tokens as independent training samples is very inefficient. Instead, we “assemble” the sampled tokens back into the sentences they belong to and implement label masks to only train on the sampled tokens in each sentence. Following this method, we can train on sampled tokens with minimal overheads.
Training and Generating New Weak Teachers. For each iteration, $t$, of EnsemW2S-AdaBoost we train a new weak teacher model $h^{t}_{\theta}$ on the sampled data, $S^{\prime}$.
Incorporating Prior Term. Following Hastie et al. (2009), multi-class boosting uses an additional $\log(c-1)$ term, where $c$ is the number of classes, in the calculation of the AdaBoost parameter $\alpha$. This term serves two purposes: (1) It enables the generation of weak models with accuracy between 50% and random $\frac{1}{c}$, which is crucial for smaller models and challenging tasks that cannot achieve 50% accuracy. (2) It ensures that $\alpha$ remains positive. Bayesian inference is used to provide proof of the benefits of this prior term. Given the large vocabulary size in our case, we introduce a prior term $\log(\frac{1}{1-\epsilon_{pre}}-1)$, where $\epsilon_{pre}$ is the pre-trained model error of the chosen LLM. This term is sensible because it represents the error before fine-tuning the LLM, effectively replacing the random error baseline. Thus, the final $\alpha$ equation is: $\alpha_{t}\leftarrow\log({\frac{1-\epsilon_{t}}{\epsilon_{t}}})+\log(\frac{1}{1-\epsilon_{pre}}-1)$.
Algorithm 1 Main Algorithm: EnsemW2S-AdaBoost
@@
1:An “easy” Q/A training dataset with $m$ examples: $S^{e}=\{(\bm{x}^{e}_{i},\bm{y}^{e}_{i})\}_{i=1}^{m}$; a pre-trained weak teacher model $h_{\theta}^{0}$ parameterized by $\theta$; total number of EnsemW2S-AdaBoost iterations $T$; a “hard” unlabeled (questions only) dataset with $O$ examples: $S^{h}=\{\bm{x}^{h}_{o}\}_{o=1}^{O}$
2:Weak-to-Strong Student Model $f_{\phi}(\cdot)$
@@
3:Initialize Token-Sample Weights: $D_{1}(i,j)\leftarrow\frac{1}{n}$ for all $i\in[m],j\in[k_{i}]$, where $k_{i}$ is the token length in the $i^{\text{th}}$ easy example (i.e., $\bm{y}^{e}_{i}=(\bm{y}_{i}^{e,1},\bm{y}_{i}^{e,2}...\bm{y}_{i}^{e,k_{i}})$) and $n=\sum_{i=1}^{m}k_{i}$
@@
4:Calculate pre-training error of $h_{\theta}^{0}$: $\epsilon_{pre}\leftarrow\sum_{i=1}^{m}\sum_{j=1}^{k_{i}}\mathbbm{1}\{h_{\theta}^{0}(\bm{x}^{e}_{i},\bm{y}_{i}^{e,j-1})\neq\bm{y}_{i}^{e,j}\}D_{1}(i,j)$
5:for $t\leftarrow 1$ to $T$  do
6:     Sample $S^{\prime}=\{(\bm{x^{\prime}}^{e}_{i},\bm{y^{\prime}}^{e}_{i})\}_{i=1}^{m}\text{ from }S\text{ using token-sample weights }D_{t}(i,j)$
7:     Train a new weak teacher $h_{\theta}^{t}$ on $S^{\prime}$
8:     Calculate $\epsilon_{t}=\sum_{i=1}^{m}\sum_{j=1}^{k_{i}}\mathbbm{1}\{h_{\theta}^{t}(\bm{x}^{e}_{i},\bm{y}_{i}^{e,j-1})\neq\bm{y}_{i}^{e,j}\}D_{t}(i,j)$
9:     if $\epsilon_{t}\geq\epsilon_{pre}$ then
10:         break
11:     Calculate $\alpha_{t}\leftarrow\log{\frac{1-\epsilon_{t}}{\epsilon_{t}}}+\log(\frac{1}{1-\epsilon_{pre}}-1)$

### v2

# EnsemW2S: Enhancing Weak-to-Strong Generalization with Large Language Model Ensembles

@@ Abstract
With Large Language Models (LLMs) rapidly approaching and potentially surpassing human-level performance, it has become imperative to develop approaches capable of effectively supervising and enhancing these powerful models using smaller, human-level models exposed to only human-level data. We address this critical weak-to-strong (W2S) generalization challenge by proposing a novel method aimed at improving weak experts, by training on the same limited human-level data, enabling them to generalize to complex, super-human-level tasks. Our approach, called EnsemW2S, employs a token-level ensemble strategy that iteratively combines multiple weak experts, systematically addressing the shortcomings identified in preceding iterations. By continuously refining these weak models, we significantly enhance their collective ability to supervise stronger student models. We extensively evaluate the generalization performance of both the ensemble of weak experts and the subsequent strong student model across in-distribution (ID) and out-of-distribution (OOD) datasets. For OOD, we specifically introduce question difficulty as an additional dimension for defining distributional shifts. Our empirical results demonstrate notable improvements, achieving 4%, and 3.2% improvements on ID datasets and, upto 6% and 2.28% on OOD datasets for experts and student models respectively, underscoring the effectiveness of our proposed method in advancing W2S generalization.
@@ 1 Introduction
Rapid advancements in Large Language Models (LLMs) have largely been driven by extensive internet-scale data and improvements in computational hardware. However, these vast datasets have largely reached their saturation point, offering diminishing returns in terms of new knowledge, and hardware-driven performance gains are also diminishing. Despite these limitations, the need for LLMs to effectively generalize to out-of-distribution, challenging, and unseen data remains crucial. This capability becomes even more essential as models approach and potentially exceed superhuman performance Burns et al. (2023), necessitating super-human level supervision from human-level data and models. In this paper, we focus on enhancing the generalization performance of human-level experts using limited human-level labeled data and smaller-scale models. We subsequently leverage these improved models to supervise and enhance larger-scale models on superhuman-level data. Our framework is inspired from the “superalignment” or “w2s generalization” problem by Burns et al. (2023), but extends it by also focusing on generalization to superhuman-level data using only human-level labeled data.
W2S generalization has been previously studied Sun et al. (2024); Burns et al. (2023); Ji et al. (2024); Charikar et al. (2024); Lang et al. (2024); however, several key challenges remain unresolved: (C1) Generalization to Out-of-distribution (OOD) and super-human level data. One key aspect of W2S generalization is the ability to extrapolate from human-level performance to super-human tasks using only easily accessible human-level data. While “easy” examples are abundant and cheap to label, “hard” or super-human examples often lie beyond the reach of annotators, either due to cognitive limits or prohibitive costs. In fact, for many tasks, groundtruth labels at the superhuman level may be fundamentally unavailable—raising a critical obstacle to benchmarking generalization at that level. Therefore, we believe that easy-to-hard (E2H) generalization is central to the W2S paradigm. To address this, we explicitly partition data into “easy” and “hard” categories and develop methods that promote generalization to harder questions even when trained solely on easy examples. (C2) Limitation of Single Weak Supervisors. Previous research predominantly utilizes a single weak supervisor, often leading to biased, inconsistent, or incomplete supervisory signals. In practice, multiple complementary weak experts (e.g., humans specializing in different domains) can be readily accessed. Our work leverages an ensemble approach that aggregates these diverse weak supervisors, providing a richer, more balanced supervisory signal and significantly enhancing overall model generalization. (C3) Lack of Iterative Weak Model Enhancement. Existing approaches typically treat weak models as static entities, neglecting the potential for iterative improvements through limited labeled data. Although methods such as voting (Anonymous, 2025) and debate (Du et al., 2023) have attempted model combination during inference, they require a stronger supervisor (reward model or judge) to select optimal tokens or answers. In contrast, our approach iteratively refines weak models without relying on stronger supervisory signals or additional data, thereby fostering a self-reinforcing cycle of continuous improvement. This dynamic enhancement not only promotes continuous improvement of weak models for improved w2s generalization but also holds potential for answering previously unsolved and challenging questions.
Figure 1: This figure illustrates the full pipeline of EnsemW2S. The top part shows how weak experts are iteratively improved by training on the same human-level labeled data. The bottom part depicts how these experts are combined at the token level during inference to supervise the strong student model. Training data is sampled using token-level weights, which define a distribution over training tokens. Based on the token-level errors of each trained weak model, we compute model weights ($\alpha_{t}$), which are then used during inference to compute a weighted combination of token-level probabilities. This aggregated distribution is used to generate answer tokens for previously unseen, superhuman-level questions, providing supervision to the strong model—following the setup in Burns et al. (2023). EnsemW2S reflects a realistic scenario where weak experts perform well on easy questions and are leveraged to supervise a stronger model on harder (and potentially unsolved) tasks.
To address the above challenges, we propose EnsemW2S, a novel token-level ensemble algorithm for autoregressive LLMs. Inspired by the spirit of AdaBoost—an ensemble method for binary classifiers known for improving generalization while resisting overfitting—EnsemW2S trains and combines complementary weak models at the token level. Specifically, it adjusts token probabilities in a manner analogous to controlled decoding methods (Mudgal et al., 2023b), but extends them to a multi-model setting. The approach iteratively improves newly introduced weak models based on identified shortcomings of prior models without requiring any additional labeled data.
We evaluate generalization rigorously on unseen data, distinguishing between in-distribution (ID) and OOD scenarios defined either by different datasets but the same task or question difficulty levels. We refer to this as easy-to-hard (E2H) generalization, a proxy for the unattainable goal of generalizing from human-level to superhuman performance—where groundtruth labels are inherently unavailable. By leveraging a small held-out labeled set for calibration, EnsemW2S enables effective supervision that improves both weak and strong models. As a result, the final strong model attains significantly better generalization and can even match or outperform models trained on full groundtruth datasets.
@@
(1) EnsemW2S: A novel token-level ensemble method to improve generalization to difficult data and further improve w2s generalization. (a) EnsemW2S iteratively improves weak models without using additional data and aggregates them via a probabilistic voting mechanism. This self-reinforcing cycle of improvement significantly enhances generalization. (b) We also provide theoretical justification for exponential decay in training error bounds as more weak models are incorporated (assuming our new weak learning condition holds). We then discuss how training data margin distribution improves, leading to better generalization while preventing overfitting. (c) We achieve improved generalization to unseen data by upto 4% for ID and 4.34%, 6% for OOD (easy-hard) with test data from the same dataset and a different dataset, respectively.
(2) Better weak models further improve W2S Generalization by providing more effective supervision. We achieve performance improvement of up to 3.2% on ID and 2.42%, 2.28% on OOD (easy-hard) with test data from the same dataset and a different dataset, respectively.
2 Framework
Figure 2: Overview of the EnsemW2S. EnsemW2S leverages hard examples with pseudo-labels generated by weak experts to train a strong student model. To evaluate its effectiveness, we measure Performance Gap Recovered (PGR) (Burns et al., 2023)
W2S Generalization with E2H Generalization
As a proxy for human and superhuman-level data, we split data into easy and hard, using the former to generalize to the latter. This E2H framework is a more pragmatic setting to study the (im)possibility of w2s generalization. in this framework, weak models (or human-level models) trained on simpler tasks guide a strong model (or superhuman-level model) toward solving more complex challenges, mirroring real-world conditions of limited human supervision. Figure 2 outlines EnsemW2S, where weak expert models ($h_{\theta}$) proficient in ‘easy tasks’ $(\bm{x}^{e},\bm{y}^{e})$ supervise a strong model ($f_{\phi}$) to solve (unsolvable) hard problems ($\bm{x}^{h}$). The strong student model under EnsemW2S is trained using unlabeled hard data alongside pseudo-labels, $(\bm{x}^{h},h_{\theta}(\bm{x}^{h}))$, where $h_{\theta}(\bm{x}^{h})=\text{func}(\{h_{\theta}^{t}\}_{t=1}^{T},\bm{x}^{h})$ denotes the ensemble-generated labels from weak experts. Ultimately, our goal is to improve the Performance Gap Recovered (PGR) which compares EnsemW2S against two baselines: a lower bound—weak experts trained on easy data and evaluated on hard data—and an upper bound—the performance of a strong model trained directly on hard data. The upper bound or oracle, is the strong model ($u_{\phi}$), which matches the student model in size and caliber but is trained using ground-truth labeled hard data $(\bm{x}^{h},\bm{y}^{h})$, typically unattainable in real-world scenarios. This oracle sets the upper performance limit for W2S models.
3 EnsemW2S: Improving W2S via Improved Expert Ensembles
In this section, we introduce EnsemW2S, a token-level ensemble method for autoregressive generative models such as LLMs. Figure 1 outlines the method. As classification is a subset of generation, we first discuss ensemble methods for binary classification LLMs, followed by the challenges of applying such methods to generation tasks, and finally present our proposed solution.
3.1 Motivation: AdaBoost for Binary Classification with LLMs
AdaBoost is a well-known ensemble learning algorithm designed for binary classification. We apply AdaBoost to a binary classification LLM to study W2S generalization. Historically, AdaBoost demonstrates strong generalization on unseen data by increasing the margins of training examples, which correlates with lower generalization error, even after achieving zero training error-making it ideal for this study. AdaBoost iteratively reweights training examples, and trains weak experts one at a time focusing on hard-to-classify samples (check Line 5 of Algorithm 2). The essential requirement is that each weak learner must perform slightly better than random guessing, thereby satisfying the weak learning condition. The outputs of these weak experts are aggregated through a weighted majority vote to produce pseudo-labels: $\mathbbm{1}\left(\sum_{t=1}^{T}\alpha_{t}h_{\theta}^{t}(\bm{x}^{h})>0\right)\in{0,1}$, where $\alpha_{t}$ are hyperparameters weighing each weak expert based on accuracy. For a detailed mathematical summary, refer to Appendix D. We then use this ensemble to generate answers for challenging (“hard”) questions, $\bm{x}^{h}$.
3.2 Challenges of Ensemble for Generation Task
Popular ensemble methods like AdaBoost Freund & Schapire (1997) leverage the “wisdom of the crowd” to build stronger classifiers by combining weaker ones, typically through classification scores. Inspired by AdaBoost’s ability to increase margin and improve generalization, we explore whether its principles can be extended to generation tasks. However, applying AdaBoost to autoregressive LLMs is infeasible due to several key challenges: (1) Single-Class vs Multi-Class: Unlike binary classification, each token prediction in LLMs is a multi-class problem over a large vocabulary (typically >4000 classes). Aggregating such high-dimensional probability distributions introduces significant computational overhead and potential numerical instability. (2) Autoregressive Generation: Unlike classification, where scores are combined over a fixed output space, generation tasks involve sequential prediction, where early token errors can propagate and affect future outputs. This makes ensemble aggregation over generations inherently more complex. (3) Variable-Length Outputs: LLMs generate variable-length outputs, making it non-trivial to combine their responses consistently.
3.3 EnsemW2S: LLM Token-Level Ensemble Method
To address the above challenges, we propose a multi-class, generation based, AdaBoost inspired, LLM ensemble algorithm where the number of classes corresponds to the vocabulary size. To develop this algorithm, we treat each token as an independent sample and include error for each token. Algorithm 1 provides the steps for EnsemW2S algorithm for Multiple-Choice Q/A tasks. In the sections below we provide detail of each step along with theoretical proofs. Appendix Fig. 5 gives details on algorithmic and data flow. Note, AdaBoost is a special case of EnsemW2S for binary classification task (see Algorithm 2 in Appendix D). Refer Figure 1 for outline and intuition behind our method.
@@ 3.3 EnsemW2S: LLM Token-Level Ensemble Method
Token-Level Weighting. Our method generates weights for each token within a sentence sample. We define the initial token-sample weights vector $D_{1}(i,j)\leftarrow\frac{1}{n}$ for all $i\in[m],j\in[k_{i}]$, where $n=\sum_{i=1}^{m}k_{i}$, $k_{i}$ is the number of tokens in the answer part of each sample $i$, $m$ is the total number of training data samples and $j$ is the $j^{\text{th}}$ token in a particular chosen $i^{\text{th}}$ sample. We update these weights, $D_{t}(i,j)$, for each iteration $t$ of EnsemW2S. Note here that unlike AdaBoost, EnsemW2S assigns weights to all tokens in a sample, adjusting for varying lengths, whereas AdaBoost maintains only one weight for each sample.
@@
Token-Level Data Sampling. We sample $S^{\prime}=\{(\bm{x^{\prime}}^{e}_{i},\bm{y^{\prime}}^{e}_{i})\}_{i=1}^{m}$ from $S$ using token-sample weights $D_{t}(i,j)$. By sampling with respect to probability masses $D_{t}(i,j)$ with repetition, we obtain a set of $n=\sum_{i=1}^{m}k_{i}$ tokens to train on. However, treating these $n$ sampled tokens as independent training samples is very inefficient. Instead, we “assemble” the sampled tokens back into the sentences they belong to and implement label masking to only train on the sampled tokens in each sentence. Following this novel treatment proposed, we can train on sampled tokens with minimal overheads.
Training and Generating New Weak Experts. For each iteration, $t$, of EnsemW2S algorithm we train a new weak expert model $h^{t}_{\theta}$ on the sampled data, $S^{\prime}$. In this way, EnsemW2S iteratively improves weak models and recruits additional weak models throughout the training process—the first approach to dynamically refine weak models while training the strong model. This self-reinforcing cycle of improvement significantly enhances generalization.
Algorithm 1 Main Algorithm: EnsemW2S
@@
0:  “Easy” Q/A dataset $S^{e}=\{(x_{i}^{e},y_{i}^{e})\}_{i=1}^{m}$, pre-trained weak expert model $h_{\theta}^{0}$, iterations $T$, “Hard” unlabeled (questions only) dataset $S^{h}=\{x_{o}^{h}\}_{o=1}^{O}$
0:  Weak-to-Strong Student Model $f_{\phi}(\cdot)$
@@
1:  Initialize Token-Sample Weights: $D_{1}(i,j)\leftarrow\frac{1}{n},\quad\forall i\in[m],j\in[k_{i}]$, where $k_{i}$ is token length of $\bm{y}^{e}_{i}=(\bm{y}_{i}^{e,1},\bm{y}_{i}^{e,2}...\bm{y}_{i}^{e,k_{i}})$ and $n=\sum_{i=1}^{m}k_{i}$
@@
2:  Compute Pre-Training Error of $h_{\theta}^{0}$: $\epsilon_{\text{pre}}\leftarrow\sum_{i=1}^{m}\sum_{j=1}^{k_{i}}\mathbbm{1}\{h_{\theta}^{0}(x_{i}^{e},y_{i}^{e,j-1})\neq y_{i}^{e,j}\}D_{1}(i,j)$
3:  for $t\leftarrow 1$ to $T$ do
4:     Sample $S^{\prime}=\{(x_{i}^{e},y_{i}^{e^{\prime}})\}_{i=1}^{m}$ from $S^{e}$ using $D_{t}(i,j)$
5:     Train new weak expert model $h_{\theta}^{t}$ on $S^{\prime}$
6:     Compute error: $\epsilon_{t}\leftarrow\sum_{i=1}^{m}\sum_{j=1}^{k_{i}}\mathbbm{1}\{h_{\theta}^{t}(x_{i}^{e},y_{i}^{e,j-1})\neq y_{i}^{e,j}\}D_{t}(i,j)$
7:     if $\epsilon_{t}\geq\epsilon_{\text{pre}}$ then
8:        Break
9:     Compute model weight: $\alpha_{t}\leftarrow\log\frac{1-\epsilon_{t}}{\epsilon_{t}}+\log\left(\frac{1}{1-\epsilon_{\text{pre}}}-1\right)$